OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−4, 8) and (−3, −6)

OpenStudy (anonymous):

The formula for slope is m=change in y/change in x. Plug in the values. m=(8-(-6))/(-4-(-3))

OpenStudy (anonymous):

Remember this formula for points, x1 = -4 x2 = -3 y1 = 8 y2 = -6 Slope = (y1 - y2) / (x1 - x2) = (8-(-6)) / (-4-(-3)) = 14/(-1) thus the slope is -14.